You borrow 10,000 from a loan company. The company wants you to pay it back in equal monthly installments in 10 years under a monthly interest rate of i%.

If the equivalent future amount of your payments at the end of year 10 is 33,000, what is the nominal annual rate?
Answer #1

First of all we are required to calculate the nominal interest rate. Then, we can determine the monthly installment using the calculated interest rate.

Present value = $ 10,000

Future value = $ 33,000

Time = 10 years

Let us assume the nominal rate of return = i% per year

\large FV = PV(F/P,i\%,n)

\large \implies (F/P,i\%,n) = \frac{FV}{PV}

\large \implies (1 +i)^n= \frac{FV}{PV}

\large \implies 1 +i= [\frac{FV}{PV}]^{\frac{1}{n}}

\large \implies i= [\frac{FV}{PV}]^{\frac{1}{n}} - 1

Now plug in all the values in the above equation we get

\large \implies i= [\frac{33,000}{10,000}]^{\frac{1}{10}} - 1

\large \implies i= 1.126811 - 1

\large \implies i= 0.126811

\large \implies i \approx 12.68\% \ per \ year

Nominal interest rate = 12.68% per year

Now calculating the monthly installment using nominal interest rate = 12.68% per year

Borrowed amount = $ 10,000

Time = 10 years = 120 months

\large A = 10,000(A/P,\frac{12.68}{12}\%,120)

\large \implies A = 10,000 \times \frac{\frac{0.1268}{12}}{1-(1+\frac{0.1268}{12})^{-120}}

\large \implies A = 10,000 \times 0.0147429

\large \implies A \approx \$ \ 147.43

Monthly installment = $ 147.43

Kindly contact through comment if having any query will be obliged to you for your generous support. Your help mean a lot to me, Please help me. Thank you.

Please help me!!!